FM radio stations with high signal strength in New Orleans break down into the followong categories:
Hip hop-2
Religious-6
Country-2
Rock-3
Talk radio-2
Other-5
If someone driving in NO tuned in to am FM radio station with high signal strength what would be the chances he would not get a country or hip hop station? Give anser as a fraction in lowest terms.

Answer by stanbon(57377)

FM radio stations with high signal strength in New Orleans break down into the followong categories:
Hip hop-2
Religious-6
Country-2
Rock-3
Talk radio-2
Other-5
If someone driving in NO tuned in to am FM radio station with high signal strength what would be the chances he would not get a country or hip hop station? Give anser as a fraction in lowest terms.
---------------------
Ways to succeed: 4
Total stations: 20
---
P(country or hip hop) = 4/20 = 1/5
====================================
Cheers,
Stan H.